Complete the missing (_92^238)U + (_6^12)C → ______ + 4 (_0^1)n
1
Expert's answer
2019-11-21T06:26:27-0500

Atomic number of U=92

Atomic number of C=6

Atomic number of missing element = 92+6-(0×4)=98

mass number of U =238

mass number of C= 12

mass number of n= 1

Mass number of missing element = 238+12-(1×4)=246

The missing element is (98^246)Cf.

An isotope of Californium.(Answer)